A wire carries a current I_1 = 5.0 A, a second wir
Answer #1

The magnetic field due to a straight wire carrying current i is
B = \mu0 i / 2 \pi r
Where r is the distance from the wire to the point where the magnetic field is measured
The magnetic field due to the wire carrying the current I1 at the point p is
B1 =  \mu0 I1 / 2 \pi r
Where r = sqrt (202  + 102)
r = 22.36 cm = 0.2236 m
B1 = 4 \pix 10-7 x 5 / 2 \pix 0.2236
B1 = 44.72 x 10-7 T
The magnetic field due to the wire carrying the current I2 at the point P is
B2  = \mu0 I2 / 2 \pi r
Where r = 10 cm = 0.1 m
B2  = 4 \pix 10-7 x 5 / 2 \pix 0.1 = 100 x 10-7 T
These two magnetic fields are oriented in different directions
The angle between these two fields should be found ,from the figure
tan \theta  = 20 cm / 10 cm = 2
\theta  = 63.430
The angle between the fields is \theta  = 1800 - 63.430  = 116.570
The resultant field is
B = sqrt (B12  + B22  + 2 B1 B2 cos \theta)
B = sqrt ( 44.722 + 1002  + 2 x 44.72 x 100 x cos 116.570)
B = 89.51 x 10-7 T